OpenStudy (anonymous):

-4-(1-5)-(-4)^2 Kinda lost someone work this out step

OpenStudy (anonymous):

-4-(1-5)-(-4)^2=-4-1+5-16=-16

OpenStudy (anonymous):

-4-(-4)-(16) -4+4-16 =-16 You want to do the exponents first then the numbers in parentheses... PEMDAS.

OpenStudy (anonymous):

okay thank you!!<33